The following preferred stocks are returning 8.5% to their owners who purchased the shares when they were issued. Calculate the prices at which they were issued. Round your answers to the nearest dollar.

Stock Dividend Current Price
A 5% $15
B 6 41
C 10 130
  1. Stock A.
    $   
  2. Stock B.
    $   
  3. Stock C.
    $   
Answer #1

Issue price= current return × current price / dividend percentage

A= 8.5% × 15 /5%= 0.085× 300

A= 25.5

B= 8.5% × 41/6%= 0.085 ×683.33

B= 58.08

C=8.5% × 130/10%=0.085×

C= 110.5